# G-Adaptivity: optimised graph-based mesh relocation for finite element methods

### v2

@@ 1 Introduction
Finite element methods (FEM) are currently the most widely-used tool for the large scale solution of partial differential equations (PDEs) (Ainsworth & Oden, 1997; Cotter, 2023). Central advantages are robustness, reliable error estimates, and thoroughly developed code bases (such as deal.II (Africa et al., 2024), DUNE (Bastian et al., 2008), Fenics (Logg et al., 2012), and Firedrake (Ham et al., 2023)), which are highly parallelisable and efficient. However, even with such optimised software the simulation of large scale problems (e.g. weather forecasting, structural simulations in engineering systems) is computationally costly. An important ingredient that determines the cost is the number of degrees of freedom (DOFs) required by a FEM to satisfy a chosen error tolerance. Since this cost depends on the number $N_{z}$ of mesh points ${\mathbf{z}}^{(i)}$ of the underlying computational mesh, it is desirable to keep $N_{z}$ moderate. Mesh adaptivity based on mesh refinement and/or relocation to capture important solution features at the right scale can balance computational cost and accuracy. However, classical mesh-adaptive methods can be difficult to implement and require significant computational resources. In contrast, in the present work, we introduce a cheap, stable and highly efficient graph neural network (GNN) architecture to implement learnable mesh relocation ($r$-adaptivivity). Our method keeps $N_{z}$ fixed and adapts the mesh point locations to reduce the overall FE error. Many classical mesh relocation methods have focused on finding and minimising mathematical substitutes of the FE error (usually simplified upper bounds) and solving additional (differential) equations to relocate the mesh points. For example, one can solve the Monge-Ampère (MA) equation (Budd et al., 2009) to find mesh point locations that minimise the interpolation error, which is an upper bound (up to some parameters and constants) of the FEM error arising from Céa’s lemma (Huang & Russell, 2011). Recent Machine Learning (ML) approaches rely on similar mathematical simplifications and learn a surrogate for the mesh equations (Song et al., 2022; Zhang et al., 2024) leading to significant speed-up with comparable error reduction. In the present work we take an entirely different approach. We present G-adaptivity, an approach to mesh adaptivity that trains a GNN to generate meshes that directly minimise the error of the corresponding FEM solution. We couple backpropagation through a novel diffusion-based GNN-deformer, with mesh point gradients obtained through an application of Firedrake adjoint (Mitusch et al., 2019; Ham et al., 2023), to minimise the FEM approximation error directly (as opposed to the upper bound considered in (Huang & Russell, 2011)). The result is a model capable of outperforming the current state-of-the-art $r$-adaptive methods, whilst retaining the significant acceleration of ML based approaches (cf. Figure 1).
@@ 3.1 Problem specification
where $\alpha,\beta\in\mathbb{R}$. For transient problems, $\Omega=\widetilde{\Omega}\times(0,T)$, where $\widetilde{\Omega}$ is a $(d-1)$-dimensional spatial domain and $(0,T)$ is the time-interval of interest. In this case, we employ the method of lines and combine the FEM with suitable time-stepping schemes (Hairer & Wanner, 1996). To compute finite element solutions, we introduce a mesh $\mathcal{T}$ of the spatial domain with $N_{z}$ nodes, which we collect in the node set $\mathcal{Z}$. The mesh $\mathcal{T}$ is used to construct trial and test functions with local support to discretize (1). For example, for a Poisson problem with homogeneous Dirichlet boundary conditions we consider the space of piecewise linear functions $S_{\mathcal{Z}}$ on $\mathcal{T}$ and solve: Find $U_{\mathcal{Z}}\in S_{\mathcal{Z}}$ such that
@@ 3.2 Adaptive Meshing
Relocation based $r$-adaptivity turns a mesh with a certain topology into another mesh with the same topology. For this, the mesh points ${\mathbf{z}}^{(i)}$ are moved, but their connectivity (and hence the associated data structures) is unaltered. Such methods typically map a fixed mesh in a computational domain (i.e. a representation of the mesh graph) to a deformed mesh in the physical domain where the PDE is posed. We denote the mesh points in the physical domain by $\mathcal{Z}=\{{\mathbf{z}}^{(i)}\}_{i=1}^{N_{z}}$ which form a triangulation $\mathcal{T}$ of $\Omega$ so that
$\mathcal{T}$ =$=\left\{\Delta^{(j)}\subset\mathcal{Z};|\Delta^{(j)}|=d+1\right\}_{j=1}^{N_{\mathcal{T}}}.$
We define the following domains and coordinates: $\Omega_{C}\subseteq\mathbb{R}^{d}$ is the “computational” domain $\Omega_{C}$ is mapped to the “physical” domain $\Omega_{P}\subseteq\mathbb{R}^{d}$ and coordinates $\boldsymbol{\xi}\in\Omega_{C}$ in the computational domain are mapped to coordinates $\boldsymbol{z}\in\Omega_{P}$ in the physical domain. To construct an adaptive mesh we consider a differentiable, possibly time-dependent, deformation map $\mathbf{F}:\Omega_{C}\rightarrow\Omega_{P}$, so that $\boldsymbol{z}=\mathbf{F}(\boldsymbol{\xi},t)$ and $\mathbf{F}{(\partial\Omega_{C})=}\partial\Omega_{P}$. If ${\boldsymbol{\xi}}^{(i)}$ are the fixed mesh points in the computational domain then $\boldsymbol{z}^{(i)}=\mathbf{F}(\boldsymbol{\xi}^{(i)},t)$. Assuming the mesh in the computational domain is regular, then determining the (properties of the) mesh in the physical domain, reduces to finding, (and analysing), ${\mathbf{F}}$.
@@ 4.1 Graph-based adaptive mesh refinement
A mesh $\mathcal{T}$ (i.e. a triangulation of the domain $\Omega$) with meshpoints $\mathcal{Z}$ gives rise to a natural graph, with the nodeset $\mathcal{V}=\mathcal{Z}$ and the edgeset =$\mathcal{E}=\{(\mathbf{z}^{i},{\mathbf{z}}^{j})\in\mathcal{V}\times\mathcal{V};\exists T\,\in\mathcal{T},\text{\ s.t.\ }{\mathbf{z}}^{i},{\mathbf{z}}^{j}\in T\}$, i.e. two nodes share an edge if there is a triangle in the mesh $\mathcal{T}$ which has both nodes as vertices. The graph $(\mathcal{V},\mathcal{E})$ can be enriched with node features $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d_{0}}:i\in\mathcal{V}\right\}$ represented in matrix notation as $\mathbf{X}\in\mathbb{R}^{N_{z}\times d_{0}}$. For example we could associate to each mesh point $\mathbf{z}^{(i)}$ the value of the solution field $u(\mathbf{z}^{(i)})$ as a feature. Likewise, we can introduce latent features that propagate through repeated application of a map on the graph, this is used in our architecture (cf. Figure 2). Mesh connectivity is stored in the adjacency matrix $\mathbf{A}$ (where $a_{ij}=1$ if $(i,j)\in\mathcal{E}$ and zero otherwise). We denote by $\mathbf{D}$ the degree matrix, the diagonal matrix whose entries $d_{i}$ count the number of nodes connected to the $i$-th node . The (random walk) graph Laplacian $\Delta$ is the operator acting on features over $\mathcal{V}$ defined as $(\Delta\mathbf{x})_{i}=x_{i}-\sum_{j\sim i}x_{j}/d_{i}$. In matrix form, $\Delta=\mathbf{I}-\mathbf{D^{-1}}\mathbf{A}$.
A graph neural network (GNN) $\mathcal{M}_{\theta}:\mathbb{R}^{N_{z}\times d_{0}}\times\mathcal{E}\rightarrow\mathbb{R}^{N_{z}\times d_{N}}$ is a map from features to features constructed with layers $\mathcal{L}_{\theta_{k}}:\mathbb{R}^{N_{z}\times d_{k}}\rightarrow\mathbb{R}^{N_{z}\times d_{k+1}}$ acting node wise as
@@
where $\varphi_{\theta_{k}}$ is the learnable edge-wise operation and $\phi_{\theta_{k}}$ a learnable node-wise aggregation.
@@
Integral to the G-adaptivity framework is the construction of the feature matrix such that the GNN can act as a mesh deformer. Similar to (Zhang et al., 2024) we construct the feature matrix by concatenating coordinates of a regular mesh $\boldsymbol{\xi}\in\mathbb{R}^{N_{x}\times d}$ with a learnable feature encoding $\mathbf{h}_{\theta}(\mathcal{Z}_{0},H)$ which, motivated by (4), is dependent on the Frobenius norm of the Hessian $H(U_{\mathcal{Z}^{0}})=\|\partial_{i}\partial_{j}u\|_{F}:\Omega\rightarrow\mathbb{R}$ of the FEM solution $U_{\mathcal{Z}^{0}}$ on the undeformed mesh $\mathcal{Z}^{0}$. When higher order finite element functions are used in the approximation space of $U_{\mathcal{Z}}$ the Hessian can be obtained by simple differentiation, but even in the case of linear elements this information is recoverable using widely-used techniques such as the one described in Appendix A.2. The final input feature matrix is then $\mathbf{X}=\left(\mathcal{Z}^{0}\mathbin{\|}\mathbf{X}_{\lambda}^{0}\right)\in\mathbb{R}^{N_{x}\times(d+|\lambda|)},\,\mathbf{X}_{\lambda}^{0}=\mathbf{h}_{\theta}(\mathcal{Z}^{0},H),$ which is passed into a GNN mesh deformer that then outputs the relocated mesh points. Previous works (Song et al., 2022; Zhang et al., 2024) used a graph attention network (GAT) (Veličković et al., 2018) as the GNN mesh deformer
where $\mathbf{X}=(\mathcal{Z},\mathbf{X}_{\lambda})$, i.e. to prevent mesh crossing the non-linearity and channel mixing are excluded from the positional channels. In the above $\mathbf{A}_{\theta}(\mathbf{X}^{k})$ is row-stochastic meaning that the top row of (6) corresponds to a graph-based averaging over graph neighbours. Motivated by (Chamberlain et al., 2021b) and velocity-based methods for meshpoint relocation introduced in section 3.2, in our G-Adaptive framework this average is replaced by a diffusion based deformer (henceforth referred to as Diffformer)
which is solved to a finite time $\tau$ and leads to the meshpoint update $\mathcal{Z}^{k+1}=\mathcal{Z}(\tau)$, i.e. an overall deformer of the form
$\begin{pmatrix}\mathcal{Z}^{k+1}\\ \mathbf{X}_{\lambda}^{k+1}\end{pmatrix}=\begin{pmatrix}\mathcal{Z}(\tau)\\ \sigma_{\lambda}\left(\mathbf{A}_{\theta}(\mathbf{X}^{k})\mathbf{X}_{\lambda}^{k}\mathbf{W}_{\lambda}\right)\end{pmatrix}.$ (8)
@@ 4.2 Structural regularization
The architectural changes between (6) and (7) lead to several regularity properties that we refer to as structural regularization. The Diffformer based architecture has a key advantage over other velocity based methods in the generation of regular meshes. A requirement of FEM meshes is that they are not ‘tangled’, i.e. that they form a well-posed triangulation of the domain $\Omega$ (i.e. no triangles overlap). This follows if each mesh point is in the convex hull of its neighbours on the graph and can equivalently be charaterised using the Jacobian of the deformation map $\mathcal{M}$.
@@
If the diffusion equation (7) is solved with the forward Euler method, then for sufficiently small pseudo-timestep $d\tau$, the discrete mesh evolution under the deformation map $\mathcal{M}$ preserves element orientations, ensuring that no mesh tangling occurs.
@@ 4.3 Firedrake adjoint optimal gradient computation and direct FEM loss
Training the GNN $\mathcal{M}_{\theta}$ requires computing the derivative of $E(\mathcal{Z})=E(\mathcal{Z},U_{\mathcal{Z}})$ with respect to the node coordinates $\mathcal{Z}$. Since evaluating $E(\mathcal{Z},U_{\mathcal{Z}})$ requires solving the PDE (1) first, a naïve application of automated differentiation would result in the solution of additional $(N_{z}\times d)$ PDEs ($N_{z}\times(d-1)$ in the transient case). To avoid the additional computational cost, we employ the well-established method of adjoints. Specifically, we employ Firedrake’s automated adjoint capabilities implemented in pyadjoint (Mitusch et al., 2019). With pyadjoint, derivatives with respect to mesh coordinates can be computed in an automated fashion as shape derivatives of $E(\mathcal{Z},U_{\mathcal{Z}})$ in directions discretized with vector-valued linear Lagrangian FEs (Ham et al., 2019).
@@
along each finite element direction $V\in S_{\mathcal{Z}}^{d}$ ($V\in S_{\mathcal{Z}}^{(d-1)}$ in the transient case). This is necessary because the shape derivative of a FE function in a direction discretized with FEs is zero (Ham et al., 2019).

### v3

@@ 1 Introduction
Finite element methods (FEM) are currently the most widely-used tool for the large scale solution of partial differential equations (PDEs) (Ainsworth & Oden, 1997; Cotter, 2023). Central advantages are robustness, reliable error estimates, and thoroughly developed code bases (such as deal.II (Africa et al., 2024), DUNE (Bastian et al., 2008), Fenics (Logg et al., 2012), and Firedrake (Ham et al., 2023)), which are highly parallelisable and efficient. However, even with such optimised software the simulation of large scale problems (e.g. weather forecasting, structural simulations in engineering systems) is computationally costly. An important ingredient that determines the cost is the number of degrees of freedom (DOFs) required by a FEM to satisfy a chosen error tolerance. Since this cost depends on the number $N_{z}$ of mesh points ${\mathbf{z}}^{(i)}$ of the underlying computational mesh, it is desirable to keep $N_{z}$ moderate. Mesh adaptivity based on mesh refinement and/or relocation to capture important solution features at the right scale can balance computational cost and accuracy. However, classical mesh-adaptive methods can be difficult to implement and require significant computational resources. In contrast, in the present work, we introduce a cheap, stable and highly efficient graph neural network (GNN) architecture to implement learnable mesh relocation ($r$-adaptivity). Our method keeps $N_{z}$ fixed and adapts the mesh point locations to reduce the overall FE error. Many classical mesh relocation methods have focused on finding and minimising mathematical substitutes of the FE error (usually simplified upper bounds) and solving additional (differential) equations to relocate the mesh points. For example, one can solve the Monge-Ampère (MA) equation (Budd et al., 2009) to find mesh point locations that minimise the interpolation error, which is an upper bound (up to some parameters and constants) of the FEM error arising from Céa’s lemma (Huang & Russell, 2011). Recent Machine Learning (ML) approaches rely on similar mathematical simplifications and learn a surrogate for the mesh equations (Song et al., 2022; Zhang et al., 2024) leading to significant speed-up with comparable error reduction. In the present work we take an entirely different approach. We present G-adaptivity, an approach to mesh adaptivity that trains a GNN to generate meshes that directly minimise the error of the corresponding FEM solution. We couple backpropagation through a novel diffusion-based GNN-deformer, with mesh point gradients obtained through an application of Firedrake adjoint (Mitusch et al., 2019; Ham et al., 2023), to minimise the FEM approximation error directly (as opposed to the upper bound considered in (Huang & Russell, 2011)). The result is a model capable of outperforming the current state-of-the-art $r$-adaptive methods, whilst retaining the significant acceleration of ML based approaches (cf. Figure 1).
@@ 3.1 Problem specification
where $\alpha,\beta\in\mathbb{R}$. For transient problems, $\Omega=\widetilde{\Omega}\times(0,T)$, where $\widetilde{\Omega}$ is a $(d-1)$-dimensional spatial domain and $(0,T)$ is the time-interval of interest. In this case, we employ the method of lines and combine the FEM with suitable time-stepping schemes (Hairer & Wanner, 1996). To compute finite element solutions, we introduce a mesh $\mathcal{T}$ of the spatial domain with $N_{z}$ nodes, which we collect in the node set $\mathcal{Z}$. The mesh $\mathcal{T}$ is used to construct trial and test functions with local support to discretize (1). For example, for a Poisson problem with homogeneous Dirichlet boundary conditions we consider the space of piecewise linear functions (vanishing on $\partial\Omega$) $S_{\mathcal{Z}}$ on $\mathcal{T}$ and solve: Find $U_{\mathcal{Z}}\in S_{\mathcal{Z}}$ such that
@@ 3.2 Adaptive Meshing
Relocation based $r$-adaptivity turns a mesh with a certain topology into another mesh with the same topology. For this, the mesh points ${\mathbf{z}}^{(i)}$ are moved, but their connectivity (and hence the associated data structures) is unaltered. Such methods typically map a fixed mesh in a computational domain (i.e. a representation of the mesh graph) to a deformed mesh in the physical domain where the PDE is posed. Note that, for transient problems, $\Omega$ should be replaced by $\tilde{\Omega}$ and $d$ should be replaced by $d-1$ in the following explanation (cf. Section 3.1). We denote the mesh points in the physical domain by $\mathcal{Z}=\{{\mathbf{z}}^{(i)}\}_{i=1}^{N_{z}}$ which form a triangulation $\mathcal{T}$ of $\Omega$ with $N_{\mathcal{T}}$ mesh elements, i.e.
$\mathcal{T}$$\subset\left\{\Delta^{(j)}\subset\mathcal{Z};|\Delta^{(j)}|=d+1\right\},\quad|\mathcal{T}|=N_{\mathcal{T}}.$
We define the following domains and coordinates: the “computational” domain $\Omega_{C}$ is mapped to the “physical” domain $\Omega_{P}\subseteq\mathbb{R}^{d}$, and the “computational” coordinates $\boldsymbol{\xi}\in\Omega_{C}$ are mapped to the “physical” coordinates $\boldsymbol{z}\in\Omega_{P}$. To construct an adaptive mesh we consider a differentiable, possibly time-dependent, deformation map $\mathbf{F}:\Omega_{C}\rightarrow\Omega_{P}$, so that $\boldsymbol{z}=\mathbf{F}(\boldsymbol{\xi},t)$ and $\mathbf{F}(\partial\Omega_{C})=\partial\Omega_{P}$. If ${\boldsymbol{\xi}}^{(i)}$ are the fixed mesh points in the computational domain then $\boldsymbol{z}^{(i)}=\mathbf{F}(\boldsymbol{\xi}^{(i)},t)$. Assuming the mesh in the computational domain is regular, then determining the (properties of the) mesh in the physical domain, reduces to finding, (and analysing), ${\mathbf{F}}$.
@@ 4.1 Graph-based adaptive mesh refinement
For simplicity of exposition we focus our discussion on the 2D case, but note that this approach generalises in a straightforward manner to 3D cases as shown in Section 5.5. A mesh $\mathcal{T}$ (i.e. a triangulation of the domain $\Omega$) with meshpoints $\mathcal{Z}$ gives rise to a natural graph, with the nodeset $\mathcal{V}=\mathcal{Z}$ and the edgeset =$\mathcal{E}=\{(\mathbf{z}^{i},{\mathbf{z}}^{j})\in\mathcal{V}\times\mathcal{V};\exists\Delta\,\in\mathcal{T},\text{\ s.t.\ }{\mathbf{z}}^{i},{\mathbf{z}}^{j}\in\Delta\}$, i.e. two nodes share an edge if there is a triangle in the mesh $\mathcal{T}$ which has both nodes as vertices. The graph $(\mathcal{V},\mathcal{E})$ can be enriched with node features $\left\{\mathbf{x}_{i}\in\mathbb{R}^{d_{0}}:i\in\mathcal{V}\right\}$ represented in matrix notation as $\mathbf{X}\in\mathbb{R}^{N_{z}\times d_{0}}$. For example we could associate to each mesh point $\mathbf{z}^{(i)}$ the value of the solution field $u(\mathbf{z}^{(i)})$ as a feature. Likewise, we can introduce latent features that propagate through repeated application of a map on the graph, this is used in our architecture (cf. Figure 2). Mesh connectivity is stored in the adjacency matrix $\mathbf{A}$ (where $a_{ij}=1$ if $(i,j)\in\mathcal{E}$ and zero otherwise). A graph neural network (GNN) $\mathcal{M}_{\theta}:\mathbb{R}^{N_{z}\times d_{0}}\times\mathcal{E}\rightarrow\mathbb{R}^{N_{z}\times d_{N}}$ is a map from features to features constructed with layers $\mathcal{L}_{\theta_{k}}:\mathbb{R}^{N_{z}\times d_{k}}\rightarrow\mathbb{R}^{N_{z}\times d_{k+1}}$ acting node wise as
@@
where $\varphi_{\theta_{k}}$ is the learnable edge-wise operation, $\phi_{\theta_{k}}$ is a learnable node-wise aggregation and $\mathcal{N}_{i}=\{j\in\mathcal{V};(i,j)\in\mathcal{E}\}$ is the set of nodes adjacent to the meshpoint $i$.
Integral to the G-adaptivity framework is the construction of the feature matrix such that the GNN can act as a mesh deformer. Similar to (Zhang et al., 2024) we construct the feature matrix by concatenating coordinates of a regular mesh $\boldsymbol{\xi}\in\mathbb{R}^{N_{z}\times d}$ with a learnable feature encoding $\mathbf{h}_{\theta}(\mathcal{Z}_{0},H)$ which, motivated by (4), is dependent on the Frobenius norm of the Hessian $H(U_{\mathcal{Z}^{0}})=\|\partial_{i}\partial_{j}u\|_{F}:\Omega\rightarrow\mathbb{R}$ of the FEM solution $U_{\mathcal{Z}^{0}}$ on the undeformed mesh $\mathcal{Z}^{0}$. When higher order finite element functions are used in the approximation space of $U_{\mathcal{Z}}$ the Hessian can be obtained by simple differentiation, but even in the case of linear elements this information is recoverable using widely-used techniques such as the one described in Appendix A.2. The final input feature matrix is then $\mathbf{X}=\left(\mathcal{Z}^{0}\mathbin{\|}\mathbf{X}_{\lambda}^{0}\right)\in\mathbb{R}^{N_{x}\times(d+|\lambda|)},\,\mathbf{X}_{\lambda}^{0}=\mathbf{h}_{\theta}(\mathcal{Z}^{0},H),$ where $\lambda$ denotes the index set for the node features, which is passed into a GNN mesh deformer that then outputs the relocated mesh points. Previous works (Song et al., 2022; Zhang et al., 2024) used a graph attention network (GAT) (Veličković et al., 2018) as the GNN mesh deformer
where $\mathbf{X}=(\mathcal{Z},\mathbf{X}_{\lambda})$ and $\mathbf{W}_{\lambda}$ is a learnable linear transformation matrix. To prevent mesh crossing the non-linearity and channel mixing are excluded from the positional channels in (6). In the above $\mathbf{A}_{\theta}(\mathbf{X}^{k})$ is row-stochastic meaning that the top row of (6) corresponds to a graph-based averaging over graph neighbours. Motivated by (Chamberlain et al., 2021b) and velocity-based methods for meshpoint relocation introduced in section 3.2, in our G-Adaptive framework this average is replaced by a diffusion based deformer (henceforth referred to as Diffformer)
which is solved to a finite end time $\tau=\tau_{end}$ and leads to the meshpoint update $\mathcal{Z}^{k+1}=\mathcal{Z}(\tau_{end})$, i.e. an overall deformer of the form
$\begin{pmatrix}\mathcal{Z}^{k+1}\\ \mathbf{X}_{\lambda}^{k+1}\end{pmatrix}=\begin{pmatrix}\mathcal{Z}(\tau_{end})\\ \sigma_{\lambda}\!\left(\mathbf{A}_{\theta}(\mathbf{X}^{k})\mathbf{X}_{\lambda}^{k}\mathbf{W}_{\lambda}\right)\end{pmatrix}.$ (8)
@@ 4.2 Structural regularization
The architectural changes between (6) and (8) lead to several regularity properties that we refer to as structural regularization. The Diffformer based architecture has a key advantage over other velocity based methods in the generation of regular meshes. A requirement of FEM meshes is that they are not ‘tangled’, i.e. that they form a well-posed triangulation of the domain $\Omega$ (i.e. no triangles overlap). This follows if each mesh point is in the interior of the convex hull of its neighbours on the graph and can equivalently be characterised using the Jacobian of the deformation map $\mathcal{M}$.
@@
If the diffusion equation (7) is solved with the forward Euler method, then for sufficiently small pseudo-timestep $d\tau<1/2$, the discrete mesh evolution under the deformation map $\mathcal{M}$ preserves element orientations, ensuring that no mesh tangling occurs.
@@ 4.3 Firedrake adjoint optimal gradient computation and direct FEM loss
Training the GNN $\mathcal{M}_{\theta}$ requires computing the derivative of $E(\mathcal{Z})=E(\mathcal{Z},U_{\mathcal{Z}})$ with respect to the node coordinates $\mathcal{Z}$. Since evaluating $E(\mathcal{Z},U_{\mathcal{Z}})$ requires solving the PDE (1) first, a naïve application of automated differentiation would result in the solution of additional $(N_{z}\times d)$ PDEs ($N_{z}\times(d-1)$ in the transient case). To avoid the additional computational cost, we employ the well-established method of adjoints. Specifically, we employ Firedrake’s automated adjoint capabilities implemented in pyadjoint (Mitusch et al., 2019). With pyadjoint, derivatives with respect to mesh coordinates can be computed in an automated fashion as shape derivatives of $E(\mathcal{Z},U_{\mathcal{Z}})$ in directions discretized with vector-valued linear Lagrangian FEs (Ham et al., 2019a).
@@
along each finite element direction $V\in S_{\mathcal{Z}}^{d}$ ($V\in S_{\mathcal{Z}}^{(d-1)}$ in the transient case). This is necessary because the shape derivative of a FE function in a direction discretized with FEs is zero (Ham et al., 2019a).